### v1

# DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation

DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Environmental perception systems play a critical role in high-precision mapping and autonomous navigation, with LiDAR serving as a core sensor that provides accurate 3D point cloud data. How to efficiently process unstructured point clouds while extracting structured semantic information remains a significant challenge, and in recent years, numerous pseudo-image-based representation methods have emerged to achieve a balance between efficiency and performance. However, they often overlook the structural and semantic details of point clouds, resulting in limited feature fusion and discriminability. In this work, we propose DAGLFNet, a pseudo-image-based semantic segmentation framework designed to extract discriminative features. First, the Global-Local Feature Fusion Encoding module is used to enhance the correlation among local features within a set and capture global contextual information. Second, the Multi-Branch Feature Extraction network is employed to capture more neighborhood information and enhance the discriminability of contour features. Finally, a Feature Fusion via Deep Feature-guided Attention mechanism is introduced to improve the precision of cross-channel feature fusion. Experimental evaluations show that DAGLFNet achieves 69.83% and 78.65% on the validation sets of SemanticKITTI and nuScenes, respectively. The method balances high performance with real-time capability, demonstrating great potential for LiDAR-based real-time applications.
1 Introduction
Semantic segmentation has emerged as a cornerstone technology for three-dimensional environmental perception, enabling dense semantic annotation and feature learning from LiDAR point clouds [1, 2, 3]. By providing structured interpretations of complex environments, this capability is fundamental to applications such as robotics and autonomous driving. Modern LiDAR sensors capture tens of millions of points per second, allowing an unprecedentedly detailed representation of the surrounding 3D structure [4, 5]. The central challenge, however, lies in devising effective strategies to process inherently unstructured and unordered point cloud data in order to extract discriminative features for reliable perception.
Current LiDAR segmentation techniques adopt several distinct paradigms for handling point cloud data. Point-based methods process raw point sets directly, enabling dense feature interaction but relying on computationally intensive neighborhood searches to capture local geometric structures [6, 7, 8]. This leads to considerable resource demands and constrains their scalability in large outdoor environments. Voxel-based methods discretize point clouds into regular volumetric grids and apply sparse convolutions for feature extraction [9, 10, 11, 12]. Yet, the cubic growth of memory consumption with resolution makes high-resolution voxel grids prohibitively costly to construct and process. Hybrid strategies that integrate features from multiple representational domains have shown improved predictive accuracy [13]. Nevertheless, the heavy computational burden and issues of system robustness continue to pose major challenges for real-world deployment.
Recently, range-image-based point cloud semantic segmentation has attracted increasing attention as it offers a favorable balance between computational efficiency and segmentation accuracy [3, 14, 15, 16]. By projecting irregular and complex LiDAR point clouds into structured two-dimensional representations such as range images [17, 18], this approach introduces several notable advantages. The regular image grid greatly simplifies neighborhood queries and local feature aggregation, while circumventing costly and time-consuming three-dimensional operations, and achieves promising performance, as shown in the Figure 1(b). However, this projection inevitably introduces structural distortions, causing blurred boundaries and semantic ambiguities that undermine the fidelity of structural representations, as shown in Figure 1(a). The principal challenge, therefore, lies in compensating for the loss of three-dimensional geometric information during projection, particularly in distant and inherently sparse regions.
(a) Average pooling of encoded features.
(b) Semantic segmentation result.
Figure 1: (a) Visualization of feature ambiguity and boundary blurring through average pooling of encoded feature channels from the LiDAR point cloud representation, and (b) corresponding semantic segmentation result demonstrating the classification performance.
A fundamental challenge of pseudo-image-based methods lies in the partitioning of point clouds, where points from distinct semantic categories may be assigned to the same subregion. Such inconsistent grouping undermines the stability of local feature encoding, causing the pseudo-image to lose critical information regarding intra-set geometric and semantic relationships, and thereby diminishing the expressive value of the segmented regions. Furthermore, when features of these local point sets are projected into a two-dimensional image, high-dimensional geometric details are compressed into abstract 2D representations, leading to blurred boundaries and loss of fine-grained structural information, which exacerbates semantic ambiguity and feature degradation. In addition, the coarse fusion of 2D image features with the original point cloud representations can introduce redundant or conflicting information, limiting the discriminative capacity and fidelity of point-wise features. These issues are particularly pronounced in distant, sparsely populated, or occluded regions, where projected subregions often contain insufficient valid points, resulting in a marked decline in the structural fidelity of the pseudo-image representation.
To address the aforementioned challenges, we propose a novel framework, DAGLFNet, which integrates global–local feature aggregation, multi-branch feature extraction, and an attention-driven fusion mechanism. This design achieves a favorable balance between accuracy and computational efficiency in pseudo-image-based point cloud semantic segmentation. Specifically, local point sets generated through sub-region partitioning based on azimuth and laser-beam segmentation often exhibit significant internal geometric variability, which undermines the stability of local feature representations. To address this challenge and enhance the consistency of local geometric structures, we propose a Global-Local Feature Fusion Encoding (GL-FFE) module, capable of simultaneously capturing global contextual dependencies and fine-grained local geometric relationships. During the mapping of point cloud subsets into image representations, boundary features are often blurred and subject to interference from adjacent regions. To overcome this limitation, we propose a Multi-Branch Feature Extraction network (MB-FE), designed to expand the receptive field and enhance the discriminative representation of boundary features. In addition, we introduce a Feature Fusion via Deep Feature-guided Attention (FFDFA) strategy during the feature integration stage, which explicitly leverages distance information as a weighting constraint to enhance the precision of cross-channel feature fusion.
In summary, our main contributions are as follows:
  1. 1.
    We introduce a novel network architecture, DAGLFNet, for semantic segmentation of LiDAR point clouds. Within this framework, geometric features of the point cloud are tightly integrated with two-dimensional pseudo-image representations, enabling efficient processing of unstructured and unordered point data while fully capturing discriminative point-wise features.
  2. 2.
    We propose a comprehensive feature enhancement strategy, comprising three key modules: (i) a GL-FFE module to capture long-range dependencies and stabilize local geometric representations; (ii) a MB-FE network to expand the receptive field and strengthen boundary feature expression; and (iii) a FFDFA mechanism that leverages distance-aware weighting to improve inter-channel feature integration.
  3. 3.
    Extensive experiments on two widely adopted LiDAR segmentation benchmarks demonstrate the superiority of DAGLFNet. Specifically, DAGLFNet achieves mean Intersection-over-Union (mIoU) scores of 69.8% and 78.7% on the validation sets of SemanticKITTI and nuScenes, respectively. Moreover, the framework can be successfully deployed on embedded platforms, enabling real-time semantic segmentation.
2 Related Work
2.1 LiDAR Point cloud segmentation
Point cloud semantic segmentation represents a cornerstone task in 3D perception [19, 20], aiming to assign precise semantic labels to individual points within a scene. By transforming raw and unstructured LiDAR data into dense, semantically enriched representations, this process enables a structured and fine-grained understanding of complex environments [21]. Among the diverse strategies developed for LiDAR point cloud semantic segmentation, point-based, voxel-based, and multi-modal data fusion approaches have emerged as three representative paradigms.
Point-based methods operate directly on raw point clouds, preserving the full spatial geometric information of the scene. PointNet [8] pioneered the application of multi-layer perceptrons (MLPs) [22] to extract global feature correlations, inspiring subsequent architectures that integrate point convolution [23, 24] to capture local spatial patterns, graph convolution [25, 26] to model neighborhood relationships, and attention mechanisms [27, 28] to focus on salient features. These methods improved the ability of point cloud networks in local feature extraction and geometric structure preservation. However, while point convolution enhanced local pattern recognition, graph convolution improved neighborhood modeling, and attention mechanisms highlighted important features, all these approaches still suffered from high computational complexity when processing large-scale point clouds due to their dense operations and lack of efficient downsampling strategies. While RandLA-Net [6] mitigates computational demands through random sampling, point-based approaches remain challenged by the efficient processing of large-scale point clouds, limiting their applicability in real-time scenarios.
To impose structure on inherently unorganized point cloud data, voxel-based approaches discretize the point cloud into regular three-dimensional grids and employ 3D convolutional neural networks (3D CNNs) to extract hierarchical geometric features. Nevertheless, the intrinsic sparsity of point clouds renders voxel grids largely empty. OctNet [29] employs hierarchical octree decomposition to handle sparse point clouds, but still faces computational limitations in large-scale scenes. MinkNet [12] leverages Minkowski convolution operations to better capture geometric relationships in sparse voxel grids. Cylinder3D [11] introduces cylindrical voxelization specifically designed for LiDAR data, improving feature representation in cylindrical coordinate systems. However, these methods collectively face challenges in balancing computational efficiency with detailed feature extraction, as the voxel count in extensive LiDAR scenes still increases rapidly, incurring significant memory and computational overhead that limits their applicability in real-time scenarios.
Recognizing the limitations of single data sources in fully capturing point cloud information, researchers have developed multi-modal fusion strategies to exploit the complementary strengths of diverse inputs. UniSeg [13] fuses voxel, view, and image features to fully utilize multi-modal semantic information. However, in practical applications, the heterogeneity of data sources in coordinate systems, resolution, and temporal synchronization makes precise feature alignment a significant challenge.
2.2 Semantic Segmentation Based on Range Images
Recently, researchers have proposed a range of methodologies to optimize range-image-based semantic segmentation, aiming to enhance its precision [30, 14, 15, 16]. RangeNet++ [30] projects LiDAR point clouds onto spherical range images and applies convolutional neural networks for semantic segmentation, effectively reformulating the task into a structured two-dimensional representation. SalsaNext [31] and CENet [17] leverage enhanced contextual modeling to effectively mitigate boundary blurring and detail degradation inherent to pseudo-image-based methodologies. RangeFormer [16] partitions the entire LiDAR scan into multiple view-specific subsets and incorporates a Transformer-based architecture to capture long-range dependencies. Although these methods demonstrate reliable overall performance, they operate solely on points projected onto the image, struggle to resolve conflicts arising from multiple points mapping to the same location, overlook occluded points, and fail to preserve the full three-dimensional structure. To address this problem, FRNet [14] integrates point cloud and 2D features, preserving the fidelity of three-dimensional geometric information while leveraging the computational efficiency of two-dimensional convolutions. FARVNet [15] incorporates a reflectivity reconstruction module, amplifying the contribution of reflectivity to feature representation and thereby enhancing model robustness. Although achieving strong overall performance, these methods overlook intra-subset feature correlations, limiting their ability to accurately recognize sparsely occluded regions at long distances.
Distinct from prior approaches, this work emphasizes the coherent modeling of intra-subset feature relationships while systematically accounting for the effects of spatial distance, highlighting a more holistic consideration of local and contextual dependencies.
Figure 2: The proposed DAGLFNet framework consists of key components such as GL-FFE, MB-FE, and FFDFA, which are responsible for contextual and geometric feature extraction, boundary enhancement, and multi-scale feature integration, respectively. Multiple stacked DAGLFNet units continuously learn complex hierarchical features from the point cloud. The Fusion Head combines point-level and group-level features to predict the final output.
3 Methodology
To overcome the limitations of pseudo-image-based segmentation, including point projection conflicts and incomplete 3D structure preservation, we propose a unified framework that effectively integrates global and local features. Our framework consists of three complementary components: a GL-FF module for capturing contextual and local geometric dependencies, a MB-FE network for enhancing boundary representations, and an AD-FF strategy for precise integration of multi-scale information. This design balances computational efficiency with high-fidelity 3D feature representation, addressing the key challenges inherent to range-image-based point cloud segmentation. The architecture overview of the proposed DAGLFNet framework is depicted in Figure 2.
3.1 Problem Definition
Given a point cloud $P={\{p_{1},p_{2},..,p_{N}\}}$ acquired by the LiDAR sensor, our network aims to assign a unique semantic label $L$ to each point, taking the point coordinates and reflectivity $p=(x,y,z,I)$ as input. This process can be summarized as:
$L=\mathcal{G}(P,\theta)$ (1)
where $\theta$ represents the learnable parameters within the network. As illustrated in Figure 2, DAGLFNet comprises four key steps: 1) a global–local point cloud feature encoder for point cloud representation; 2) an image feature encoder for extracting semantic features; 3) a point cloud feature fusion module guided by depth values via an attention mechanism; 4) a fusion head that integrates multi-level point cloud features to achieve precise semantic prediction.
3.2 Feature Encoder
Given the limited information available from point clouds, effective feature encoding is critical [32]. Unlike voxelization or regional partitioning [11, 30], we divide the point cloud into $M$ groups $P=\{\mathbb{P}_{1},\mathbb{P}_{2},...,\mathbb{P}_{M}\}$ as follows:
$r=\sqrt{x^{2}+y^{2}+(z-h_{l})^{2}},\alpha=\text{atan(y,x)},\phi=\phi_{l},$ (2)
where $h_{l}$ and $\phi_{l}$ refer to the $l$-th Velodyne sensor. We then compute the projection coordinates of points $(r,\alpha,\phi)$ as $u=((\alpha+\pi)2\pi)W$, $v=l$, where $(u,v)$ denotes the grid coordinate of a point in the range image with a resolution of $H\times W$.
We first obtain the virtual center of each group by applying average pooling over all points within the group, denoted as $p_{avg}$. Based on this, the initial features of each point comprising its 3D coordinates, reflectance, depth, and offsets relative to the $p_{avg}$, are represented as $\bar{p}_{j}\in\mathbb{R}^{1\times 10}$. The features for each point are given by:
$\bar{p}_{j}=([x,y,z,depth,I;p_{j}-p_{avg}])$ (3)
Collectively, the features of all points form the point cloud feature matrix $\bar{P}=\{\bar{p}_{1},\bar{p}_{2},...,\bar{p}_{j}\}\in{\mathbb{R}^{N\times 10}}$. Subsequently, MLP [22] is employed to extract point cloud features. Specifically, the per-point features are first encoded by an MLP, yielding per-point feature representation $\mathcal{F}_{p}^{0}$. Group-level features are then obtained by aggregating the features of all points within each group using both maximum and average pooling. The outputs of these two pooling operations are combined to form the final group-level representation. The outputs of these two pooling operations are fused and projected to yield the final group-level representation $\mathcal{F}_{g}^{0}$:
$\left\{\begin{aligned} \mathcal{F}_{p}^{0}&=\text{MLP}(\bar{P}),\\ F_{\text{cat}}&=[\,\text{MAX}(\mathcal{F}_{p}^{0});\,\text{AVE}(\mathcal{F}_{p}^{0})\,],\\ \mathcal{F}_{g}^{0}&=\text{Flat}\big(\text{ReLU}(\text{Linear}(F_{\text{cat}}))\big).\end{aligned}\right$. (4)
where $\text{Flat}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{H\times W\times C}$ denotes the function that maps point features onto the range image plane with resolution $(H,W)$. In our network, $\mathcal{F}_{p}^{0}$ and $\mathcal{F}_{g}^{0}$ serve as inputs, $\mathcal{F}_{g}^{0}$ encodes global point cloud information while integrating group-level features, forming a comprehensive representation that captures both global contextual dependencies and local geometric structures, and enhances feature expressiveness.
3.3 Image Feature Extraction
Considering that the input image features are projections of group-level features, which leads to sparsity and blurred boundary contours, we propose a multi-branch feature extraction architecture to enhance feature representation. Following prior work [14, 33, 16], we construct the backbone using multiple convolutional blocks. The input of the $i$-th stage is denoted as $F_{p}^{i-1}$ and $F_{g}^{i-1}$, which correspond to the feature map generated by the previous stage. The proposed BasicBlock is a multi-branch residual unit consisting of three parallel branches, designed to capture features at different receptive fields. Branch 1) employs a standard $3\times 3$ convolution to extract local features; Branch 2) uses a dilated $3\times 3$ convolution (dilation=2) [34] to enlarge the receptive field and capture richer contextual information; Branch 3) focuses on edge enhancement by first reducing the channel dimension with a $1\times 1$ convolution, followed by a $3\times 3$ convolution. Subsequently, the outputs of the three branches are concatenated along the channel dimension. A $1\times 1$ convolution is then applied to fuse these features into a unified representation. Finally, the input is added to the fused output through a residual connection, followed by an activation function, enabling direct information flow and enhancing feature representation. This step can be expressed in the following form:
$\left\{\begin{aligned} {F}_{1}&=\sigma\big(\text{Conv}_{3\times 3}(\mathcal{F}_{g}^{i-1}\big),\\[3.0pt] {F}_{2}&=\sigma\big(\text{Conv}_{3\times 3}^{d=2}(\mathcal{F}_{g}^{i-1})\big),\\[3.0pt] {F}_{3}&=\sigma\big(\text{Conv}_{3\times 3}\big(\sigma(\text{Conv}_{1\times 1}(\mathcal{F}_{g}^{i-1}))\big)\big),\\[3.0pt] {F}_{\text{temp}}&=\text{Conv}_{1\times 1}([{F}_{1},{F}_{2},{F}_{3}]),\\[3.0pt] {\tilde{F}_{g}^{i-1}}&=\sigma\big({F}_{\text{temp}}+{F}_{g}^{i-1}\big)\end{aligned}\right$. (5)
Here, $\text{Conv}_{k\times k}$ denotes a $k\times k$ convolution, $Conv_{3\times 3}^{d=2}$ denotes a dilated convolution with a dilation rate of 2, and $\sigma(\cdot)$ represents the normalization and activation applied after the convolution.
3.4 Feature Update Module
We take point-level and group-level features as input, and to preserve feature consistency, the local group-level features are integrated into the point-level features through a feature fusion mechanism that maintains spatial relationships and enhances discriminative power. We first project the group-level feature representation $\tilde{F}_{g}^{i-1}$ into its corresponding point-level feature space $\tilde{F}_{p}^{i-1}$ according to the projection index, using the operator $\text{InFlat}(\cdot):\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{N\times C}$. The resulting point-level features are then combined with both the ${F}_{p}^{i-1}$ and the depth features of the point cloud within a tailored attention framework. Unlike conventional attention mechanisms [35, 36], our approach introduces depth information as a dynamic modulation factor to adaptively refine feature weighting. This depth-guided adjustment preserves spatial coherence during feature transformation and effectively mitigates the geometric distortions that typically arise in cross-dimensional feature fusion:
$\left\{\begin{aligned} V_{map}&=Linear(\text{InFlat}(\tilde{F}_{g}^{i-1}))\\ V_{pts}&=Linear({F}_{p}^{i-1})\\ F_{fuse\_p}&=W_{map}\odot V_{map}+W_{pts}\odot V_{pts}\\ \mathcal{F}_{p}^{i}&=Linear(F_{fuse\_p})\end{aligned}\right$. (6)
The depth values depthdepth are linearly transformed to generate the query $Q$. The $\tilde{F}_{p}^{i-1}$ are projected into key and value representations, $K_{\text{map}}$ and $V_{\text{map}}$, respectively, while the $F_{p}^{i-1}$ are similarly transformed into $K_{\text{pts}}$ and $V_{\text{pts}}$. The attention weights $W_{\text{map}}$ and $W_{\text{pts}}$ are computed by measuring the similarity between the depth query $Q$ and the corresponding keys $K_{\text{map}}$ and $K_{\text{pts}}$, followed by softmax normalization. These weights are then applied to their respective value matrices and aggregated to obtain the fused representation $F_{{fuse\_p}}$. Finally, $F_{{fuse\_p}}$ undergoes a linear transformation to produce the output feature $\mathcal{F}_{p}^{i}$.
Given the inherent blurriness of image features, the convolutional process intrinsically degrades feature discriminability, compromising representational robustness. To counteract the degradation of image feature discriminability introduced by convolutional operations, the updated point-level features $\mathcal{F}_{p}^{i}$ are re-projected onto the image space by $\text{Flat}(\cdot)$, concatenated with $\mathcal{F}_{g}^{i-1}$, and fused through convolutional processing:
$\tilde{F}_{fuse\_g}=\sigma(\text{Conv}_{1\times 1}([\text{Flat}(\mathcal{F}_{p}^{i});\tilde{\mathcal{F}}_{g}^{i-1}]))$ (7)
To alleviate feature degradation caused by multiple convolutional operations, the fused feature $\tilde{F}_{g}$ is integrated with the global-level feature $\mathcal{F}_{fuse\_g}^{i-1}$ at the current stage through a residual-attentive enhancement mechanism:
$\mathcal{F}_{g}^{i}=\mathcal{F}_{g}^{i-1}+\phi(\text{Conv}_{3\times 3}(\sigma(\text{Conv}_{3\times 3}(\tilde{F}_{fuse\_g}))))\odot\tilde{F}_{fuse\_g}$ (8)
where $\phi(\cdot)$ represents the normalization and the sigmoid activation applied after the convolution.
3.5 Fusion Head Module
To fully leverage the fine-grained spatial details provided by low-level features and the rich semantic context encoded in high-level features for complementary advantages, the primary task of the Fusion Head is to aggregate features from multiple stages. Specifically, for point-level features, the consistent topological structure across stages allows for direct feature-level concatenation, enabling the integration of multi-stage information while preserving local details and semantic cues. However, due to the downsampling operations in the backbone network, image features from different stages exhibit varying spatial resolutions. To address this issue, all image features are resized to a consistent spatial resolution using linear interpolation to achieve spatial alignment, denoted as $\text{BilInterp}(\cdot)$:
[…]
where $K$ denotes the number of network layers, and $(h,w)$ represents the target resolution obtained through linear interpolation. We observe that the point-level feature $\mathcal{F}_{p}^{out}$ mainly contains local spatial details for describing fine structural characteristics, while the group-level feature $\mathcal{F}_{g}^{out}$ integrates semantic information from a broader receptive field to represent the overall scene characteristics.
$\mathcal{F}_{logit}=MLP(MLP(InFlat(\mathcal{F}_{g}^{out}))+\mathcal{F}_{p}^{out})+\mathcal{F}_{p}^{0}$ (10)
Here, $\mathcal{F}_{logit}$ is used to generate the final semantic scores with a linear head for the point over the entire point cloud.
4 Experiments
This section evaluates the robustness of DAGLFNet and the balance between accuracy and computational efficiency. Benchmark datasets and implementation protocols are first outlined. Quantitative and qualitative experimental results demonstrate that DAGLFNet attains an optimal trade-off between performance and efficiency, highlighting distinctive advantages. Performance and efficiency are further validated on embedded platforms, confirming practical applicability. Comprehensive ablation studies elucidate the contribution of each network component.
4.1 Datasets
We conducted comprehensive evaluations on two widely used autonomous driving LiDAR datasets. SemanticKITTI [4], collected in Karlsruhe, Germany using a Velodyne HDL-64E sensor, comprises 22 sequences with annotated point clouds, where Sequences 0–7 and 9–10 are used for training, Sequence 8 for validation, and Sequences 11–21 for online testing. Each scene contains roughly 120,000 points, annotated across 28 semantic categories, with a vertical field of view spanning $25^{\circ}$ to $3^{\circ}$. nuScenes [5], captured using a 32-beam LiDAR, includes 1,000 driving scenes with dense point clouds annotated with 32 classes; for semantic segmentation, 16 categories are evaluated. Its vertical field of view ranges from $30^{\circ}$ to $10^{\circ}$. Together, these datasets provide diverse and challenging scenarios for evaluating LiDAR-based semantic segmentation methods.
4.2 Evaluation Metrics
We use mean Intersection over Union (mIoU) and mean Accuracy (mAcc) to evaluate the semantic segmentation performance on point clouds. Suppose the dataset contains multiple semantic classes, and let $TP_{cla}$, $FP_{cla}$, and $FN_{cla}$ denote the number of true positives, false positives, and false negatives for class clacla, respectively. Then, the IoU for class clacla is defined as:
${IoU}_{cla}=\frac{TP_{cla}}{TP_{cla}+FP_{cla}+FN_{cla}}.$ (11)
The mean IoU (mIoU) across all classes is computed as:
$\text{mIoU}=\frac{1}{\text{number of classes}}\sum_{cla}\text{IoU}_{cla}.$ (12)
Similarly, the per-class accuracy is defined as:
$\text{Acc}_{cla}=\frac{TP_{cla}}{TP_{cla}+FN_{cla}},$ (13)
and the mean accuracy (mAcc) across all classes is:
$\text{mAcc}=\frac{1}{\text{number of classes}}\sum_{cla}\text{Acc}_{cla}.$ (14)
These metrics provide a comprehensive measure of segmentation quality, accounting for both overlap and class-wise correctness.
4.3 Implementation Details
Although SemanticKITTI [4] and nuScenes [5] differ in the number of LiDAR beams, both datasets share a full $360^{\circ}$ horizontal field of view. The range image resolutions are set to $64\times 1024$ and $32\times 1024$ for SemanticKITTI and nuScenes, respectively. The network depth is configured as [3, 4, 6, 3] to balance representational capacity and computational efficiency. The optimization strategy employs the AdamW [37] optimizer with an initial learning rate of 0.001, while the OneCycle [38] policy is adopted to adaptively adjust the learning rate throughout the training process. The batch size is fixed at 4 to maintain a trade-off between efficiency and memory consumption. Unless otherwise specified, all experiments were conducted on a single NVIDIA RTX 4090 GPU.
4.4 Quantitative Results
We systematically compare DAGLFNet with several state-of-the-art network architectures to comprehensively evaluate performance across multiple benchmark datasets. The evaluation encompasses both accuracy and computational efficiency, demonstrating that DAGLFNet achieves an excellent balance between the two and exhibits outstanding overall capability.
Table 1: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [4] val set. All mIoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU Car Bicycle Motorcycle Truck Other-vehicle Person Bicyclist Motorcyclist Road Parking Sidewalk Other-ground Building Fence Vegetation Trunk Terrain Pole Traffic-sign RandLA-Net [6] $50.0$ $92.0$ $8.0$ $12.8$ $74.8$ $46.7$ $52.3$ $46.0$ $0.0$ $93.4$ $32.7$ $73.4$ $0.1$ $84.0$ $43.5$ $83.7$ $57.3$ $73.1$ $48.0$ $27.0$ RangeNet++ [30] $51.0$ $89.4$ $26.5$ $48.4$ $33.9$ $26.7$ $54.8$ $69.4$ $0.0$ $92.9$ $37.0$ $69.9$ $0.0$ $83.4$ $51.0$ $83.3$ $54.0$ $68.1$ $49.8$ $34.0$ SequeezeSegV2 [39] $40.8$ $82.7$ $15.1$ $22.7$ $25.6$ $26.9$ $22.9$ $44.5$ $0.0$ $92.7$ $39.7$ $70.7$ $0.1$ $71.6$ $37.0$ $74.6$ $35.8$ $68.1$ $21.8$ $22.2$ SequeezeSegV3 [40] $53.3$ $87.1$ $34.3$ $48.6$ $47.5$ $47.1$ $58.1$ $53.8$ $0.0$ $95.3$ $43.1$ $78.2$ $0.3$ $78.9$ $53.2$ $82.3$ $55.5$ $70.4$ $46.3$ $33.2$ SalasNet [31] $59.4$ $90.5$ $44.6$ $49.6$ $86.3$ $54.6$ $74.0$ $81.4$ $0.0$ $93.4$ $40.6$ $69.1$ $0.0$ $84.6$ $53.0$ $83.6$ $64.3$ $64.2$ $54.4$ $39.8$ MinkowskiNet [41] $58.5$ $95.0$ $23.9$ $50.4$ $55.3$ $45.9$ $65.6$ $82.2$ $0.0$ $94.3$ $43.7$ $76.4$ $0.0$ $87.9$ $57.6$ $87.4$ $67.7$ $71.5$ $63.5$ $43.6$ SPVNAS [42] $62.3$ $96.5$ $44.8$ $63.1$ $55.9$ $64.3$ $72.0$ $86.0$ $0.0$ $93.9$ $42.4$ $75.9$ $0.0$ $88.8$ $59.1$ $88.0$ $67.5$ $73.0$ $63.5$ $44.3$ Cylinder3D [11] $64.9$ $96.4$ $\mathbf{61.5}$ $78.2$ $66.3$ $69.8$ $80.8$ $\underline{93.3}$ $0.0$ $94.9$ $41.5$ $78.0$ $1.4$ $87.5$ $55.0$ $86.7$ $72.2$ $68.8$ $63.0$ $42.1$ PMF [43] $63.9$ $95.4$ $47.8$ $62.9$ $68.4$ $75.2$ $78.9$ $71.6$ $0.0$ $\mathbf{96.4}$ $43.5$ $80.5$ $1.0$ $88.7$ $60.1$ $\underline{88.6}$ $\underline{72.7}$ $\underline{75.3}$ $65.5$ $43.0$ rangvit [44] $60.9$ $94.7$ $44.1$ $61.4$ $71.9$ $37.7$ $65.3$ $75.5$ $0.0$ $95.5$ $48.4$ $83.1$ $0.0$ $88.3$ $60.0$ $86.3$ $65.3$ $72.7$ $63.1$ $42.7$ CENet [17] $61.5$ $91.6$ $42.4$ $61.7$ $82.4$ $63.5$ $64.4$ $76.6$ $0.0$ $93.0$ $50.3$ $72.7$ $0.1$ $85.0$ $54.4$ $84.1$ $61.0$ $70.3$ $55.2$ $42.8$ RangeFormer [16] $66.5$ $95.0$ $58.1$ $72.1$ $85.1$ $59.8$ $76.9$ $86.4$ $0.2$ $94.8$ $\underline{55.5}$ $81.7$ $\underline{13.0}$ $88.5$ $64.5$ $86.5$ $66.8$ $73.0$ $64.0$ $52.0$ SphereFormer [45] $67.8$ $96.8$ $51.0$ $75.0$ $\mathbf{93.4}$ $64.4$ $77.0$ $92.6$ $\underline{0.8}$ $94.7$ $53.2$ $52.1$ $3.7$ $90.7$ $58.5$ $\mathbf{88.7}$ $71.3$ $\mathbf{75.9}$ $64.7$ $\mathbf{54.5}$ FRNet [14] $67.6$ $97.2$ $53.3$ $72.9$ $81.5$ $72.9$ $77.2$ $90.8$ $0.2$ $95.9$ $53.7$ $83.9$ $9.0$ $90.5$ $65.9$ $87.0$ $66.8$ $72.6$ $64.0$ $47.9$ waffleIron [46] $68.0$ $96.1$ $58.1$ $\underline{79.7}$ $77.4$ $59.0$ $\underline{81.1}$ $92.2$ $\mathbf{1.3}$ $95.5$ $50.2$ $83.6$ $6.0$ $\mathbf{92.1}$ $67.5$ $87.8$ $\mathbf{73.8}$ $73.0$ $65.7$ $\underline{52.2}$ FARVNet [15] $68.5$ $97.0$ $54.2$ $75.9$ $89.6$ $72.6$ $76.0$ $90.1$ $0.0$ $95.7$ $\mathbf{56.9}$ $83.4$ $\mathbf{22.7}$ $89.8$ $62.3$ $87.0$ $65.9$ $73.0$ $63.2$ $47.1$ DAGLFNet $\underline{69.1}$ $\underline{97.4}$ $58.2$ $78.0$ $89.6$ $\underline{76.6}$ $80.5$ $92.3$ $0.0$ $96.0$ $50.1$ $\underline{83.7}$ $0.00$ $91.3$ $\underline{68.5}$ $87.9$ $69.5$ $74.1$ $\underline{66.3}$ $51.0$ DAGLFNet $\mathbf{69.9}$ $\mathbf{97.4}$ $\underline{59.9}$ $\mathbf{81.4}$ $\underline{90.9}$ $\mathbf{77.3}$ $\mathbf{81.3}$ $\mathbf{93.5}$ $0.0$ $\underline{96.2}$ $51.7$ $\mathbf{84.1}$ $0.1$ $\underline{91.8}$ $\mathbf{70.0}$ $88.0$ $70.5$ $74.2$ $\mathbf{67.3}$ $51.6$
Table 2: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [5]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU Barrier Bicycle Bus Car Construction-vehicle Motorcycle Pedestrian Traffic-cone Trailer Truck Driveable-surface Other-ground Sidewalk Terrain Manmade Vegetation AF2S3Net [47] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [30] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [48] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [49] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [31] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [50] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [44] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [11] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [51] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [52] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ WaffleIron [46] $77.6$ $78.7$ $\mathbf{51.3}$ $93.6$ $88.2$ $47.2$ $86.5$ $\underline{81.7}$ $68.9$ $69.3$ $83.1$ $96.9$ $74.3$ $75.6$ $74.2$ $87.2$ $85.2$ RangeFormer [16] $78.1$ $78.0$ $45.2$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $\mathbf{85.6}$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [45] $\underline{78.4}$ $77.7$ $43.8$ $\underline{94.5}$ $93.1$ $52.4$ $\underline{86.9}$ $81.2$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ WaffleAndRange [53] $77.6$ $78.5$ $\underline{49.6}$ $91.8$ $87.6$ $52.7$ $86.7$ $\mathbf{82.2}$ $70.1$ $67.2$ $79.7$ $97.0$ $74.7$ $\underline{76.8}$ $74.9$ $87.5$ $85.0$ FRNet [14] $76.1$ $77.2$ $39.5$ $93.4$ $88.6$ $52.1$ $81.4$ $75.1$ $65.7$ $66.2$ $79.7$ $96.8$ $75.3$ $75.4$ $75.9$ $88.4$ $85.4$ FARVNet [15] $77.8$ $77.8$ $42.1$ $\mathbf{94.5}$ $91.8$ $54.9$ $84.5$ $77.0$ $66.7$ $70.2$ $\underline{85.4}$ $97.0$ $74.6$ $76.4$ $75.9$ $89.2$ $87.4$ DAGLFNet $78.3$ $78.5$ $46.5$ $89.3$ $90.7$ $57.1$ $\mathbf{88.9}$ $79.3$ $70.3$ $69.7$ $83.1$ $97.0$ $\underline{75.9}$ $76.1$ $\underline{76.0}$ $89.9$ $88.2$ DAGLFNet $\mathbf{78.7}$ $\underline{78.8}$ $47.1$ $89.7$ $90.4$ $\underline{57.4}$ $86.8$ $80.4$ $\underline{71.1}$ $70.6$ $81.9$ $\underline{97.0}$ $\mathbf{76.5}$ $76.4$ $\mathbf{76.2}$ $90.0$ $88.4$
Tables 1 and 2 present a comparison of our method with existing state-of-the-art approaches on the SemanticKITTI [4] and nuScenes [5] validation sets. The results demonstrate that DAGLFNet significantly outperforms previous methods across most category-level metrics. Specifically, on the SemanticKITTI [4] validation set, our method achieves an improvement of 1.1 mIoU over WaffleIron [47] and 0.6 mIoU over FARVNet [15]; on the nuScenes [5] validation set, our method achieves an improvement of 0.7 mIoU over WaffleAndRange [53] and 0.5 mIoU over FARVNet [15], attaining the best or second-best performance across multiple categories.
Figure 3 further illustrates a comparison between DAGLFNet and the baseline method across randomly selected scenarios, for both dynamic and static classes. Notably, DAGLFNet achieves substantial improvements over the baseline in both types of classes, with an improvement of up to 23% observed in the truck category. Figure 4 shows the performance comparison between DAGLFNet and the baseline method across different distance ranges. DAGLFNet consistently outperforms the baseline, achieving improvements of 9.4% at 20–30m, 4.4% at 30–40m, and 12.5% at 40–45m. At closer ranges 10–20m, both methods perform similarly, while at 45–50m, DAGLFNet still provides a slight gain of 1.8%. Point cloud density progressively decreases with distance, creating a challenging environment for accurate discrimination. Remarkably, our method maintains strong effectiveness in these sparse, long-range regions.
Figure 3: Class-wise LiDAR segmentation results of DAGLFNet and the baseline model on the test set of SemanticKITTI.
Figure 4: Comparison of mIoU (%) between DAGLFNet and the baseline method across different distance ranges.
Figure 5 illustrates the relationship between mIoU and inference speed (ms/scan) on the SemanticKITTI [4] validation set for several state-of-the-art point cloud semantic segmentation methods. Our method, DAGLFNet, achieves the highest 69.1% mIoU while maintaining a moderate inference speed of 45.4 ms/scan, striking a favorable balance between accuracy and efficiency. Methods such as FRNet [14] and FARVNet [15] exhibit slightly lower accuracy with comparable inference speed, whereas faster methods like CENet [17] achieve very low latency (7.6 ms/scan) but with significantly lower accuracy (61.5% mIoU). Similarly, WaffleIron [46] shows fast inference but slightly reduced precision. Overall, Ours demonstrates a well-balanced trade-off between accuracy, inference efficiency, and model complexity, making it highly suitable for real-time applications.
Figure 5: mIoU vs. inference speed for various point cloud semantic segmentation methods on the SemanticKITTI [4] validation set. Marker size indicates model size. DAGLFNet achieves the best balance of accuracy and speed.
4.5 Qualitative Results
Figure 6 visualizes the segmentation errors in challenging scenes on the SemanticKITTI [4] validation set. Specifically, we discuss the following four scenarios: (a) for large-scale vegetation point clouds, other methods struggle to effectively extract the most discriminative features for classification, resulting in extensive segmentation errors. In contrast, DAGLFNet can accurately segment the majority of complex regions, demonstrating superior performance; (b) furthermore, in occluded regions of similar scenes, DAGLFNet can still accurately classify the vegetation within the occluded areas, demonstrating its efficiency in local feature recognition; (c) in complex intersection scenarios, where the number of categories is large and diverse, road recognition is particularly critical. However, methods such as SphereFormer [16] and WaffleIron [46] fail to correctly classify the roads, whereas DAGLFNet can accurately identify them, ensuring precise overall scene understanding; (d) the recognition of obstacles such as vehicles at intersections is equally important. However, other methods perform poorly in identifying vehicles near turning intersections, whereas DAGLFNet demonstrates more stable and reliable recognition capabilities.
Refer to captionSphereFormerwaffleironFARVNetFRNetOurs(a)(b)(c)(d)
Figure 6: Segmentation errors in challenging scenes on the SemanticKITTI [4] validation set. Red represents misclassified areas, and gray represents correctly classified areas. The four representative scenarios include: (a) segmentation of large-scale vegetation point clouds; (b) classification of vegetation in occluded regions; (c) road recognition in complex intersections; and (d) identification of vehicles and obstacles at intersections. DAGLF-Net demonstrates higher accuracy and robustness across all scenarios.
To further investigate the impact of segmentation errors, we additionally visualize three types of scenarios: (a) narrow sidewalk; (b) interference cases; and (c) multiple interferences including vegetation, terrain, and occlusions, as shown in Figure 7. Specifically, the performance in three representative scenarios is as follows: (a) in narrow sidewalk scenes, WaffleIron [46], FARVNet [15], and FRNet [14] misclassify sidewalks as terrain, whereas only DAGLFNet correctly segments them; (b) in cases where sidewalks partially occlude the road, FRNet [14] and FARVNet [15] incorrectly classify the road as sidewalk. Although WaffleIron [46] correctly segments most of the road, it still exhibits confusion in certain regions, splitting the same road area into road and sidewalk, whereas DAGLFNet effectively handles this challenging situation; (c) in regions where vegetation and terrain are interwoven, other methods fail to accurately distinguish between the two, while DAGLFNet achieves accurate segmentation with minimal errors.
Refer to captionGround TruthwaffleironFARVNetFRNetOurs(a)(b)(c)
Figure 7: Three challenging segmentation error scenarios: (a) narrow sidewalks; (b) interference cases; (c) multiple interferences including vegetation, terrain, and occlusions.
To further evaluate the performance of DAGLFNet in distant sparse regions, we compared the segmentation of vehicles and buses , as shown in Figure 8. In distant sparse point clouds, FARVNet [15] and FRNet [14] tend to confuse sparse vehicles with adjacent sparse terrain, resulting in cars being misclassified as terrain. Under distant occlusion conditions, FARVNet [15] misclassifies buses as background buildings, while FRNet [14] only partially recognizes them and still confuses them with the background. This is mainly because conventional methods do not fully exploit global context and multi-scale features when processing sparse point clouds and local information. In contrast, DAGLFNet combines local and global features, enhances contextual information, aligns multi-scale features, and integrates a depth-guided attention mechanism, enabling distant and occluded targets to be more effectively distinguished and recognized, thereby achieving higher robustness and accuracy.
Figure 8: Qualitative comparison of segmentation performance by different methods in distant, sparse, and occluded scenarios.
4.6 Ablation Study
In this section, we discuss the effectiveness of each design component in the DAGLFNet architecture. All experiments are conducted and reported separately on the validation sets of SemanticKITTI [4] and nuScenes [5].
Table 3: Ablation study of each component in Ours on the val set of SemanticKITTI [4] and nuScenes [5]. BL: BaseLine; GL-FFE: Global-Local Feature Fusion Encoding; MB-FE: Multi-Branch Feature Extraction; FFDFA: Feature Fusion via Deep Feature-guided Attention; TTA: Test Time Augmentation. All mIoU and mAcc scores are given in percentage (%).
BL GL-FFE MB-FE FFDFA TTA SemKITTI nuScenes
mIoU mAcc mIoU mAcc
$67.3$ $74.0$ $76.1$ $83.9$
$67.8$ $74.4$ $76.3$ $84.9$
$68.2$ $74.8$ $77.1$ $85.0$
$69.1$ $75.1$ $78.3$ $85.5$
$69.9$ $75.5$ $78.7$ $85.7$
By introducing GL-FFE, which integrates global and local contextual information to enrich point feature representation, the model achieves an improvement of 0.5 mIoU on SemanticKITTI [4] and 0.2 mIoU on nuScenes [5]. This indicates that capturing both global scene context and local details helps to better distinguish complex structures in sparse point clouds. Subsequently, adding MB-FE, which extracts multi-scale and diverse semantic features through parallel branches, further increases mIoU by 0.4 on SemanticKITTI [4] and 0.8 on nuScenes [5], suggesting that the multi-branch design enables more comprehensive feature extraction across varying object scales. Finally, the FFDFA module, leveraging attention guided by deep features to selectively enhance informative points while suppressing background noise, contributes an additional 0.9 mIoU on SemanticKITTI [4] and 1.2 mIoU on nuScenes [5], demonstrating its effectiveness in emphasizing critical features and improving segmentation accuracy in challenging regions. Finally, adopting test time augmentation during inference, following prior works, brings an improvement of 0.8% and 0.4% mIoU, respectively.
Range Image Representation. We investigate the effect of range image resolution on DAGLFNet performance. As the resolution decreases, the projected range images become coarser, causing a loss of fine-grained details in the point cloud representation. This reduction primarily affects the accurate extraction of local features and reduces segmentation accuracy for small or distant objects. In contrast, higher resolutions preserve more spatial details, enhance feature extraction, and improve segmentation performance, especially in sparse or complex regions. However, excessively high resolutions increase computational costs, while the performance gains remain limited. As shown in Table 5, we compare the results under different range image resolutions, and the configuration of $64\times 1024$ achieves the best balance between performance and efficiency.
Network Depth. To investigate the influence of network depth on model performance, we conduct an ablation study using different depth configurations of the proposed network, denoted as $[1,1,1,1]$, $[2,2,2,2]$, $[3,3,3,3]$, and $[3,4,6,3]$. As shown in Table 4, increasing network depth generally improves segmentation accuracy (mIoU), while slightly reducing inference speed (FPS) and increasing the number of parameters. The configuration $[3,4,6,3]$ achieves the best balance between accuracy and computational efficiency, demonstrating the effectiveness of deeper hierarchical feature extraction in our model.
Table 4: Ablation of network depth: Evaluation of different depth configurations on the SemanticKITTI [4] validation set. All mIoU scores are reported in percentage (%).
Depth mIoU FPS Params Car Bicycle Motorcycle Truck Other-vehicle Person Bicyclist Motorcyclist Road Parking Sidewalk Other-ground Building Fence Vegetation Trunk Terrain Pole Traffic-sign $[1,1,1,1]$ $66.1$ $32$ $28.2M$ $95.6$ $51.1$ $70.9$ $81.5$ $62.6$ $80.0$ $86.9$ $0.0$ $95.2$ $47.4$ $83.0$ $0.1$ $89.1$ $58.0$ $86.8$ $67.9$ $71.3$ $64.4$ $50.7$ $[2,2,2,2]$ $66.9$ $27$ $36.0M$ $96.5$ $54.0$ $70.9$ $80.5$ $64.9$ $77.61$ $89.5$ $0.1$ $95.8$ $52.0$ $83.4$ $10.4$ $89.8$ $64.0$ $86.0$ $67.6$ $71.0$ $64.7$ $50.8$ $[3,3,3,3]$ $68.3$ $24$ $43.8M$ $96.9$ $54.6$ $75.3$ $88.4$ $72.0$ $76.2$ $91.3$ $0.1$ $95.7$ $56.7$ $83.5$ $14.3$ $89.8$ $62.7$ $86.9$ $64.6$ $72.8$ $63.1$ $47.1$ $[3,4,6,3]$ $69.1$ 2121 $51.6M$ $97.4$ $58.2$ $78.0$ $89.6$ $76.6$ $80.5$ $92.3$ $0.0$ $96.0$ $50.1$ $83.7$ $0.00$ $91.3$ $68.5$ $87.9$ $69.5$ $74.1$ $66.3$ $51.0$
Table 5: Ablation of range image resolution: Evaluation of the impact of different range image resolutions on model performance using the SemanticKITTI [4] validation set.
Method mIoU FPS Car Bicycle Motorcycle Truck Other-vehicle Person Bicyclist Motorcyclist Road Parking Sidewalk Other-ground Building Fence Vegetation Trunk Terrain Pole Traffic-sign $64\times 256$ $65.8$ $30$ $95.1$ $50.0$ $67.4$ $82.7$ $55.7$ $65.3$ $87.6$ $0.0$ $95.0$ $64.7$ $83.3$ $24.7$ $88.1$ $60.6$ $88.8$ $61.2$ $79.1$ $55.8$ $45.7$ $64\times 512$ $68.2$ $26$ $96.9$ $54.3$ $75.7$ $71.3$ $77.1$ $75.1$ $91.1$ $0.0$ $95.6$ $63.6$ $84.3$ $14.3$ $90.3$ $66.5$ $88.1$ $66.7$ $75.1$ $61.6$ $48.8$ $64\times 768$ $68.0$ $23$ $96.7$ $51.7$ $79.7$ $92.2$ $69.5$ $77.4$ $85.2$ $0.0$ $95.7$ $47.3$ $83.5$ $10.0$ $90.9$ $66.3$ $87.1$ $61.6$ $72.6$ $66.2$ $50.3$ $64\times 1024$ $69.1$ 2121 $97.4$ $58.2$ $78.0$ $89.6$ $76.6$ $80.5$ $92.3$ $0.0$ $96.0$ $50.1$ $83.7$ $0.0$ $91.3$ $68.5$ $87.9$ $69.5$ $74.1$ $66.3$ $51.0$ $64\times 1280$ $68.6$ $17$ $96.2$ $54.7$ $77.2$ $82.2$ $68.4$ $79.6$ $90.0$ $0.0$ $94.9$ $51.6$ $84.0$ $0.0$ $91.3$ $68.8$ $87.6$ $68.3$ $73.1$ $64.6$ $51.7$
4.7 Failure Cases
We identify a primary limitation of our model in sparse and occluded regions when analyzing its performance in specific scenarios. As illustrated in Figure 9, when the LiDAR point cloud becomes sparse due to occlusion and contains large empty areas, DAGLFNet struggles to correctly recognize and distinguish semantically similar terrain contours and sidewalks. This occurs mainly because, in regions with insufficient point information, the model lacks sufficiently dense local geometric features to serve as reliable discriminative cues, leading to the misclassification of ambiguous boundaries and similar structures. These findings suggest that in environments with severe sparsity or occlusion, the robustness of fine-grained geometric and semantic contextual understanding still needs further enhancement.
(a) Ground Truth
(b) DAGLFNet
Figure 9: Challenging cases in sparse and occluded regions. (a) Ground Truth; (b) prediction results of DAGLFNet. In areas with sparse LiDAR points and occlusions, DAGLFNet struggles to accurately distinguish highly similar ground contours and sidewalks, highlighting its limited capability in recognizing local features when dense point cloud information is lacking.
5 Conclusion
In this work, we presented DAGLFNet, a pseudo-image-based network for point cloud semantic segmentation, designed to address challenges caused by point cloud sparsity and occlusions. By integrating local and global features, employing a multi-branch feature extraction mechanism, and leveraging deep feature-guided attention fusion, DAGLFNet achieves robust and accurate segmentation in long-range, sparse, and complex scenarios. Extensive experiments on benchmark datasets, including SemanticKITTI and nuScenes, demonstrate that the network attains competitive performance while maintaining real-time efficiency and enabling deployment on embedded devices.
We also identify a limitation of DAGLFNet in extremely sparse or occluded regions, where the model may misclassify semantically similar structures such as terrain and sidewalks, indicating that further improvement in leveraging fine-grained geometric and semantic context is needed in challenging environments.
Overall, DAGLFNet provides efficient and accurate segmentation for sparse and complex point clouds, balancing real-time performance and embedded deployment capability. Future research could focus on enhancing network understanding of geometric and semantic context in highly sparse and occluded areas, further improving robustness and accuracy in complex scenarios.
[…]
Funding Key R&D Project of the Sichuan Provincial Department of Science and Technology—Research on Three-Dimensional Multi-Resolution Intelligent Map Construction Technology (2024YFG0009), the Intelligent Identification and Assessment for Disaster Scenes: Key Technology Research and Application Demonstration (2025YFNH0008), Project of the sichuan Provincial Department ofscience and Technology—— Application and Demonstration of Intelligent Fusion Processing of Laser Imaging Radar Data (2024ZHCG0176), the "Juyuan Xingchuan" Project of Central Universities and Research Institutes in Sichuan——High-Resolution Multi-Wavelength Lidar System and Large-Scale Industry Application(2024ZHCG0190), the Sichuan Science and Technology Program, Research on Simulator Three-Dimensional View Modeling Technology and Database Matching and Upgrading Methods, and the Key Laboratory of Civil Aviation Flight Technology and Flight Safety (FZ2022KF08).